Can you please solve.

x  1  2  3  4 
a 18 16 14 12

a=????
Thanks Brad

something like x + 8 -6!!!!!!!!!!

Answer by AnlytcPhil(1806)

Can you please solve.

x  1  2  3  4 
a 18 16 14 12

a=????
Thanks Brad

something like x + 8 -6!!!!!!!!!!

-------------------------------


The successive values of "a" differ by -2
But the successive values of "x" differ by only 1
However the successive values of "2x" will differ by 2

So let's put in a list between them for 2x

 x  1  2  3  4  
2x  2  4  6  8
 a 18 16 14 12

Now we can observe that 2+18=20, 4+16=20, 6+14=20, 8+12=20

So it must be that 2x + a must equal 20, or

2x + a = 20, or solving for a

     a = 20 - 2x.

So the answer is 20 - 2x

Edwin